OpenStudy (anonymous):

the average of 5 consecutive integers starting from m as the first integer is n.What is the average of following integers? m+1,m+3,m+5,m+7,m+9

OpenStudy (anonymous):

given m+(m+1)+(m+2)+(m+3)+(m+4)=5n or 5m+10=5n or m+2=n now average of (m+1,m+3,m+5,m+7,m+9)=(m+1+m+3+m+5+m+7+m+9) /5 =(5m+25)/5 = m+5 = m+2 +3 =( m+2 )+3 =n+3 (as m+2=n)

OpenStudy (anonymous):

r u sure? @matricked

OpenStudy (anonymous):

if u know what is meant by average then i have given complete answer ... else i should tell u the average .. so do u know what is meant by average in arithmetic..

OpenStudy (anonymous):

yes i do know.

OpenStudy (anonymous):

then i hope u can well understand any mistake/error in the solution..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

the answer would be m+2?

OpenStudy (anonymous):

options are m+4 m+5 n+6 none

OpenStudy (anonymous):

@matricked

OpenStudy (raden):

it should be n+3

OpenStudy (sirm3d):

the answer is m+5 or n+3

OpenStudy (anonymous):

m+5 is also one of the answer...

OpenStudy (anonymous):

what;s the correct one?